OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center (0, 8), radius = 8

hartnn (hartnn):

Equation of circle : \(\Large (x-h)^2 + (y-k)^2 = r^2 \)

hartnn (hartnn):

with h,k as cetre, here h= 0, k=8 and r = radius = 8

OpenStudy (anonymous):

I'm supposed to fill in the blanks, they've already answered part of it.

OpenStudy (anonymous):

So here's what they've given me;

OpenStudy (anonymous):

__+(___)_______

OpenStudy (anonymous):

That's all they've said, so I have to figure out the rest =/

hartnn (hartnn):

what do you get after you plug in h=0, k= 8, r= 8 in the formula i gave you ?

OpenStudy (anonymous):

(x-0)^2+(y-8)^2=8^2

hartnn (hartnn):

yes, simplifying that x^2+ (y-8)^2 = 64 does that help ypu fill all the blanks ?

OpenStudy (anonymous):

Ohh, that does indeed. Thank you so much! :) That makes sense now!

hartnn (hartnn):

you're most welcome ^_^